From a square piece of cardboard with width x inches, a square of x - 3 inches is removed from the center. Write the area of the remaining piece as a function of x.
Answer by ptaylor(1894)

Area = length x width
Area of big square =x^2
area of small square=(x-3)^2=x^2-6x+9
Area of the remaining piece is area of big square - area of small square
Thus,

Area of remaining piece=x^2-(x^2-6x+9)
Area of remaining piece=x^2-x^2+6x-9
Area of remaining piece=6x-9
Hope this helps----ptaylor